# Learning 3D-Gaussian Simulators from RGB Videos

### v1

@@ 3.3 Dynamics transformer
Besides the computational efficiency of SPC over kNN, its main advantage lies in the compatibility with standard dot-product attention mechanisms. To understand this, we first examine how the standard PTv3 architecture computes particle-wise predictions from a single point cloud at a single time step. The process begins by extracting particle-wise embeddings $E_{i}$ for each serialized point $(s_{i},\tilde{g}_{i})$ using a sparse convolutional neural network (CNN). Next, the embedded SPCs are progressively down sampled via grid pooling before being grouped and shuffled into patch pairs. Conditional positional embeddings (xCPE) are then added to the embeddings, followed by layer normalization and a patch-wise attention layer predicting the change in the embeddings $\Delta E_{i}$. The pooling, patch shuffling, and attention blocks are arranged in a U-net [25] like architecture that first reduces the size of the SPCs in an encoding step and then mirrors this architecture. In 3DGSim, the final layer of the dynamics model predicts the change in the particle positions $\Delta p_{i}$ and the change in their features $\Delta f_{i}$.
Figure 4: The dynamics model encodes the time step into each embedding and merges embeddings from adjacent time-steps. The TEM and PTv3 blocks are applied repeatedly until all embeddings are merged.
Figure 5: Temporal point cloud serialization.
Temporal point cloud serialization
For dynamics prediction, we generate multiple latent particle clouds from different time steps $t_{1},...,t_{T}$ which are then transformed into SPCs by PTv3’s initialization block. To encode the time step $t_{k}$ into the SPC, we extent the batch-version of the serialization (4) to
[…]
with batch size $b$, the temporal integer code $s_{t_{k}}$, total time step integer $\tau=log_{2}(T)$, and $\kappa=48\,\mathrm{Bits}$. As each position dimension is encoded using 16 bits and using a grid size of $G=0.004\,\mathrm{m}$, the theoretical encoding capacity corresponds to a spatial extent of approximately $0.004\,\mathrm{m}\cdot 2^{16}\approx 216\,\mathrm{m}$ per dimension. In turn, by extending PTv3’s initialization, we obtain the temporally serialized point clouds (t-SPC) $\tilde{S}_{i}(t_{k})=\{(\tilde{s}_{i},\tilde{g}_{i})\}_{i=1}^{N_{k}}$.
Temporal encoding
As shown in Figure 6, prior to merging point clouds from different time steps, we incorporate a learned, time-step-specific positional encoding $E_{t_{k}}$ into the embeddings. Specifically, each embedding $E_{i}(t_{k})$ at time step $t_{k}$ is updated as follows:
$E_{i}(t_{k})\leftarrow E_{i}(t_{k})+E_{t_{k}}.$ (7)
This temporal encoding ensures that the attention network can distinguish points across different temporal instances, enabling the model to reason about dynamics over time. Similar positional encoding methods have previously been applied in transformer architectures to differentiate positions within sequences [39].
Temporal merging
Unlike PTv3, which limits attention computations exclusively to patches composed of points from the same time step within one batch, our method extends the receptive field across time steps. We achieve this by introducing a temporal merging operation, which applies a one-bit right shift to the temporal serialization codes $s_{t_{k}}$:
$\mathrm{Merge}(\tilde{s}_{i})=[b\mid(s_{t_{k}}>>1)\mid s_{i}].$ (8)
In our current implementation, we specifically merge pairs of consecutive time steps $t_{j}$ and $t_{j+1}$, where $j\in 2n,n\in\mathbb{N}^{*}$. We adopted this simple and computationally convenient strategy, though more sophisticated temporal merging schemes—such as merging points from larger windows of adjacent time steps—can also be explored in future work. Importantly, in contrast to [11], which require dedicated transformer modules for each time step, our temporal merging enables the reuse of the same attention block across time steps. This substantially reduces memory demands and facilitates greater knowledge transfer across time steps.
Figure 6: Temporal merging and embedding followed by patch grouping for the patch-wise attention mechanism.
3.4 Loss function
3DGSim is trained solely on an image reconstruction loss. This loss is computed from rasterized multi-view images, generated based on both the encoder predictions of past point clouds $\{\{g_{i}(t_{k})\}_{i=1}^{N_{k}}\}_{k=0}^{T}$ and the simulated future point cloud trajectory $\{\{g_{i}(t_{k})\}_{i=1}^{N_{k}}\}_{k=T+1}^{T+T^{\prime}}$. Specifically, the loss is defined as:
$\mathcal{L}=(1-\lambda)\frac{1}{T}\sum_{k=0}^{T}\mathcal{L}_{k}+\lambda\sum_{k=T+1}^{T+T^{\prime}}\gamma^{k}\mathcal{L}_{k},$ (9)
with $\lambda=0.5$, temporal decay factor $\gamma=0.87$, $T\in\{2,4\}$ and $T^{\prime}=12$. The per-frame reconstruction loss $\mathcal{L}_{k}$ measures the discrepancy between ground-truth ($I_{k}^{\text{gt}}$) and predicted ($I_{k}$) multi-view images using a weighted combination of pixel-wise $\ell_{2}$ and LPIPS [40] terms, reading
$\mathcal{L}_{k}=\alpha\,\mathcal{L}_{2}(I_{k}^{\text{gt}},I_{k})+\beta\,\mathcal{L}_{\text{LPIPS}}(I_{k}^{\text{gt}},I_{k}),$
with hyper-parameters $\alpha=1$ and $\beta=0.05$.
Figure 7: Comparison of our method to Cosmos. Cosmos is conditioned on the prompts: ”A rigid body falling on a circular gray ground.”, ”A soft body falling on a circular-gray-ground.” and ”A rigid body falling on a red rectangular cloth which is fixed on its edges.”. Metrics are computed on the same test images for all methods.
Table 2: Comparison of 3DGSim to Cosmos. ”Explicit” models use 3DGS parameters and a static latent feature as inputs to the dynamics model, while ”Latent” models use only latent features mapped to Gaussians after dynamics. The $\dagger$-model relaxes camera constraints (3 view cameras out of 6 total); the $\ddagger$-model omits segmentation masks for static elements. Metrics ”future” and ”past” are means over all timesteps. ”4-12” denotes conditioning on 4 past states with rollout prediction of 12 future steps.
Dataset Model PSNR (future) $\uparrow$ PSNR (past) $\uparrow$
rigid 3DGSim 4-12 latent 28.28 ± 2.52 32.93 ± 1.56
3DGSim 2-12 latent 28.08 ± 2.46 33.00 ± 1.62
3DGSim 4-12 explicit 27.88 ± 2.43 32.77 ± 1.57
3DGSim 2-12 explicit 27.07 ± 2.27 32.67 ± 1.65
cosmos 22.35 ± 3.82
elastic 3DGSim 4-12 latent 33.15 ± 3.51 34.55 ± 2.26
3DGSim 2-12 latent 32.05 ± 3.48 35.99 ± 1.88
3DGSim 2-12 explicit 29.92 ± 1.72 40.85 ± 2.94
3DGSim 4-12 explicit 29.69 ± 1.75 40.16 ± 3.07
3DGSim 4-12 latent $\dagger$ 31.60 ± 3.09 32.55 ± 2.12
3DGSim 4-12 latent $\ddagger$ 32.66 ± 3.43 34.45 ± 2.44
cosmos 18.87 ± 3.99
cloth 3DGSim 4-8 latent 26.98 ± 2.63 34.81 ± 2.28
3DGSim 2-8 latent 26.25 ± 2.38 35.22 ± 1.97
3DGSim 4-8 explicit 23.72 ± 1.52 39.75 ± 2.32
3DGSim 2-8 explicit 17.97 ± 2.02 35.47 ± 1.68
cosmos 21.10 ± 3.56
4 Experiments
In what follows, train 3DGSim on different datasets and test the model’s ability to generalize.
Model setup
Unless stated otherwise, the following training and parameter settings serve as defaults in the experiments. The state consists of dynamic $f^{\text{dyn}}$ and invariant features $f^{\text{inv}}$ of size $(32,32)$ for the implicit- and $(0,16)$ for the explicit 3D Gaussian particle representation. The inverse rendering encoder follows MVSplat, reducing candidate depths from 128 to 64 due to smaller scene distances. Default near-far depth ranges are [0.2, 4] for rigid bodies and [15, 8] for the other datasets. The dynamics transformer defaults to PTV3 with a 5-stage encoder (block depths [2,2,2,6,2]) and a 4-stage decoder ([2,2,2,2]). Grid pooling and temporal merging strides default to [1,4,2,2,2] and [1,2,2,2,2], respectively, with grid size $G{=}0.004$ m. Attention blocks use patches of size 1024, encoder feature dimensions [32,64,128,256,512], decoder dimensions [64,128,256], encoder heads [2,4,8,16,32], and decoder heads [4,4,8,16]. For the camera setup, we select 4 uniformly distributed views at random and an additional 5 target cameras from the remaining cameras (out of 12 total) to compute the reconstruction loss.
Datasets
The rigid body dataset consists of 1,000 simulated trajectories involving six rigid objects (turtle, sonny school bus, squirrel, basket, lacing sheep, and turboprop airplane) from the GSO dataset [41]. Each trajectory spans 32 frames at 12 FPS, providing controlled dynamics characteristic of rigid body motion. The elastic dataset, aimed at capturing plastic deformable object dynamics, includes six objects (dragon, duck, kawaii demon, pig, spot, and worm) simulated using the Genesis MPM elastoplastic simulator [15]. Each object undergoes deformation upon collision with a circular gray ground, offering scenarios of complex elastic behavior. The cloth dataset includes the same set of objects as the elastic dataset. Here, the cloth is fixed at four corners, posing the challenge to infer implicit constraints and modeling dynamic cloth-like deformations. Both elastic and cloth datasets contain 200 trajectories per object, simulated with a time step of 0.001 and 20 substeps, resulting in 2-second sequences captured at 42 FPS (84 frames per trajectory).
Metrics
For evaluation, 12% of trajectories are chosen at random and held out from each dataset, and we report each model’s PSNR (with SSIM being reported in the supplementary material). The PSNR (Peak Signal to Noise Ratio) correlates with the mean squared error between the ground truth and predicted images, and therefore captures both deviations in the dynamics and effects of blurring.
Training
Our models are trained for $\sim$120,000 steps using a cosine annealing warm-up and a learning rate of $2\times 10^{-4}$, with batch sizes of 6 and 4 for 2-step and 4-step states, respectively. To optimize memory and speed, we use gradient checkpointing and flash attention v2 [42]. Training is performed on a single H100 GPU and typically takes 5 to 6 days to complete.
4.1 Trajectory simulation
Figure 7 shows the PSNR of our model’s simulations using two or four past frames and the latent representation. To the best of our knowledge, 3DGSim is currently the only open-source, particle-based simulator trained directly on videos that integrates rendering and simulation; therefore, we compare against Cosmos-1.0-Autoregressive-5B-Video2World [3], a video generation model conditioned on 9-frame video input and language prompts. Additionally, we ablate key hyperparameters and report results in Table 2. Explicit representations achieve similar performance but generalize poorly, particularly with fewer cameras (see supplementary).
4.2 Scene editing and model generalization
Figure 8: 3DGSim’s prediction of a rigid plane captures shadows by adjusting the appearance of ground particles.
As an additional rigorous test, we simulate multi-object interactions by duplicating bodies (2 to 5 instances, see Figure 9), and run the simulation for 180 steps. Despite being trained only on object-ground collisions, 3DGSim correctly captures realistic multi-body dynamics. Instead of collapsing into chaotic interactions, individual objects retain structural integrity and move cohesively. We attribute this to the transformer architecture, which avoids structural constraints that might limit generalization, as seen in GNN approaches.
A striking consequence of removing explicit physics biases is that 3DGSim not only captures physics but also learns to reason about broader scene properties. One particularly intriguing outcome is its ability to accurately model shadows as shown in Figure 8, which requires a holistic understanding of scene geometry, lighting, and object relationships.
4.3 Simulation speed
Simulation speed is critical for robotics applications. Traditional simulators (FEM, MPM, PBD) typically employ small integration timesteps. Learned approaches enable larger timesteps, allowing 3DGSim to simulate elastic, cloth at 42 FPS and rigid dynamics at 12 FPS, with inference speeds of ${\sim}16$ FPS (4 past steps) and ${\sim}20.1$ FPS (2 past steps), using under 20 GB VRAM on an H100 GPU and achieving near real-time prediction speeds.
Figure 9: Although 3DGSim has not been trained on scenes containing this particular elastic object nor several objects, it predicts physically plausible deformations. Interestingly, the combined weight of five objects pushes the particles through the table causing color changes or falling particles.
5 Discussion
3D Gaussian simulators (3DGSim) integrate particle-based simulation with recent advancements in point-cloud transformer scalability and the fast-rendering capabilities of 3D Gaussian splatting. This combination results in an efficient and scalable simulator capable of learning complex interactions between physical particles from multi-view RGB images. By combining pooling with leveraging patch-wise attention of temporally serialized point clouds, 3DGSim effectively mitigates the inductive biases of GNNs, which rely on local message passing or global nodes.
Furthermore, recognizing the critical role of visual scene information in simulating motion, we jointly train inverse rendering alongside dynamics. As a consequence, the model does not only learn a rich visual representation of the scene but also extracts features essential for modeling dynamics such as shadows which could be informative about relative distances.
By avoiding inductive biases where possible and prioritizing scalability over first-principles, transformers have already demonstrated their effectiveness in uncovering complex patterns within large datasets. In turn, we are eager to explore 3DGSim on significantly larger datasets encompassing a diverse range of physical interactions. As we scale up, incorporating language embeddings into simulators emerges as a particularly promising direction. After all, words such as “liquid” or “mirror” could provide meaningful priors about an object’s physical properties. In this regard, we hope that 3DGSim points the way towards scalable particle-based simulators that are able to infer the interactions of complex physical phenomena from video as well as potentially text data, paving the way for robots that truly understand the dynamics of their surroundings.
Supplementary Material
Appendix A Ablations
For the ablation studies, only the elastic dataset is used. The default configuration uses the latent representation, 4-step past conditioning, 12-step future rollouts, 4 input views, 5 target views, and a total of 12 cameras for training, unless otherwise specified. Any deviations from these parameters are explicitly stated or made clear within the context of the respective ablation.
@@ Appendix A Ablations
Table 3: Rollout Length
@@
Table 4: Camera Setup (85k steps)
@@
Table 5: Segmentation Masks
@@ A.4 Modality Configurations
Table 6: Input Modality
Different input modality configurations were tested. Here, we adapt the notation ”a-b-c-d”, each varying the temporal span of input conditioning $a$ is the number of particle frames representing the state, $b$ indicates the number of backward frames used to predict $c$ future steps, and total rollout steps during training $d$, leading to $b\cdot c\cdot d$ rollout steps per training step. Both variants (”4-1-2-6”, ”4-4-1-3” attain competitive performance (PSNR of 32.59 and 32.27 respectively), significantly reducing the computational load compared to longer standard rollouts. This highlights promising avenues for future investigation, emphasizing balance between computational efficiency and performance quality.
A.5 Grid Resolution
@@
Table 7: Grid Resolution.
@@
Table 8: Temporal Merger. The models are trained for 80k steps if not specified otherwise.
We experiment with various temporal merging strides for each encoder stage combined with embedding options of timestep position encoding. Since we only train with 4 past steps, the strides ”..” don’t influence the results. After 80k iterations, results clearly indicate two critical factors for success: the use of learned positional embeddings and timing of merging operations. Optimal results occur when merging temporal information only after early spatial processing stages ([1,2,2..]), whereas early or too-late merging drastically diminishes performance. Poor outcomes with late merging likely arise due to spatial pooling operations that dilute vital temporal distinctions before merging.

### v2

@@ A.2 PTv3: Scalable point cloud transformations
Besides the computational efficiency of SPC over kNN, its main advantage lies in the compatibility with standard dot-product attention mechanisms. To understand this, we first examine how the standard PTv3 architecture computes particle-wise predictions from a single point cloud at a single time step. The process begins by extracting particle-wise embeddings $E_{i}$ for each serialized point $(s_{i},\tilde{g}_{i})$ using a sparse convolutional neural network (CNN). Next, the embedded SPCs are progressively down sampled via grid pooling before being grouped and shuffled into patch pairs. Conditional positional embeddings (xCPE) are then added to the embeddings, followed by layer normalization and a patch-wise attention layer predicting the change in the embeddings $\Delta E_{i}$. The pooling, patch shuffling, and attention blocks are arranged in a U-net Ronneberger et al. (2015) like architecture that first reduces the size of the SPCs in an encoding step and then mirrors this architecture. In 3DGSim, the final layer of the dynamics model predicts the change in the particle positions $\Delta p_{i}$ and the change in their features $\Delta f_{i}$.
This section describes additional implementation details that were not explicitely mentioned in the main paper.
Feature Encoding Network
In 4.2 we describe the feature encoding network that transforms pixel-aligned features into view-independent the latent features. To regress the FiLM conditioning we use a 2-layer CNN with GELU activation, kernel of size 3 and channel dimensions $(10,20)$. The first dimension also matches the size of the conditioning vector.
Particle Wise MLP
At the end of the dynamics model, the embedding of each particle is mapped back to the particle latent features. For that we use an MLP with shapes $(128,128)$ and GELU activation between each layer.
Appendix B Ablations
For the ablation studies of 3DGSim, only the elastic dataset is used. The default configuration uses the latent representation, 4-step past conditioning, 12-step future rollouts, 4 input views, 5 target views, and a total of 12 cameras for training, unless otherwise specified. Any deviations from these parameters are explicitly stated or made clear within the context of the respective ablation.
@@ Appendix B Ablations
Table S1: Rollout Length
@@
Table S2: Camera Setup (85k steps)
@@
Table S3: Segmentation Masks
@@ B.4 Modality Configurations
Table S4: Input Modality
Different input modality configurations were tested. Here, we adapt the notation "a-b-c-d", each varying the temporal span of input conditioning $a$ is the number of particle frames representing the state, $b$ indicates the number of backward frames used to predict $c$ future steps, and total rollout steps during training $d$, leading to $b\cdot c\cdot d$ rollout steps per training step. Both variants ("4-1-2-6", "4-4-1-3" attain competitive performance (PSNR of 32.59 and 32.27 respectively), significantly reducing the computational load compared to longer standard rollouts. This highlights promising avenues for future investigation, emphasizing balance between computational efficiency and performance quality.
B.5 Grid Resolution
@@
Table S5: Grid Resolution.
@@
Table S6: Temporal Merger. The models are trained for 80k steps if not specified otherwise.
We experiment with various temporal merging strides for each encoder stage combined with embedding options of timestep position encoding. Since we only train with 4 past steps, the strides ".." don’t influence the results. After 80k iterations, results clearly indicate two critical factors for success: the use of learned positional embeddings and timing of merging operations. Optimal results occur when merging temporal information only after early spatial processing stages ([1,2,2..]), whereas early or too-late merging drastically diminishes performance. Poor outcomes with late merging likely arise due to spatial pooling operations that dilute vital temporal distinctions before merging.
Appendix C Visualizations
In this section we show rolled out dynamic prediction for different scenarios. If not otherwise stated, the first row in each image shows the ground truth predictions. First we depict scene editability and composabilty followed by several generalizations to multi body and visualization for each dataset. Videos of the predictions below are available in the supplementary material.
C.1 Scene editability
Figure S1: Ground Removal
Figure S2: Higher Throw
Figure S3: Parallel Simulation
C.2 Generalization to multiple bodies
Figure S4: Scene with 2 bunnies.
Figure S5: Scene with 3 bunnies.
Figure S6: Scene with 5 bunnies.
Figure S7: Scene with 5 worms.
Figure S8: Various simulation scenarios with different numbers and types of objects.
C.3 Visualization of cloth simulations
Figure S9: Dragon
Figure S10: Duck
Figure S11: Cow
Figure S12: Devil
Figure S13: Various objects falling on a red cloth with its corners fixed.
C.4 Visualization for elastic dynamics
Figure S14: Cow
Figure S15: Dragon
Figure S16: Duck
Figure S17: Pig
Figure S18: Worm
Figure S19: Various elastic simulations for different objects.
C.5 Visualization for rigid body dynamics
Figure S20: Car
Figure S21: Squirrel
Figure S22: Shoe
Figure S23: Box
Figure S24: Turtle
Figure S25: Various rigid simulations for different objects.